OpenStudy (anonymous):

Becca's family has 4 gallons of milk they drink 5/8 gallon per day . how long will their milk last ? A. 6 2/5 days B. 4 days C. 2 1/2 days D. 4 5/8 days I will give medal and i will fan and write good 'bout you on your prof.

OpenStudy (anonymous):

i just need the straight uop answer btw

OpenStudy (anonymous):

Convert your gallons of milk into a fraction first

OpenStudy (anonymous):

okk

OpenStudy (anonymous):

I have answer A if thts right?

OpenStudy (anonymous):

4 / (5/8) = your answer

OpenStudy (anonymous):

i know thts not right but whatevs at least you tried :P

OpenStudy (anonymous):

Yes, A is correct

OpenStudy (anonymous):

you are correct... 4 / (5/8) = 4 * 8/5 = 32/5 = 6 2/5

OpenStudy (anonymous):

ok i will close and the title i'll put on is this Will give medal fan and write good 'bout you on youre prof

OpenStudy (anonymous):

Good job, hope I helped!

OpenStudy (anonymous):

you did alot :D